Please help with this problem solving by substitution:
3x + 4y = 8
4.5x + 6y = 12
This is what I got:
3x + 4y = 8
4y = 8 - 3x
y = 2 - .75x
Then:
3x + 4(2-.75x) = 8
3x + 8 - 3x = 8
3x - 3x = 8 - 8
0 = 0
I don't understand what I'm doing wrong. Please help!
Thank you, Alison.

Answer by stanbon(48568)

3x + 4y = 8
4.5x + 6y = 12
This is what I got:
3x + 4y = 8
4y = 8 - 3x
y = 2 - .75x
Then:
3x + 4(2-.75x) = 8
3x + 8 - 3x = 8
3x - 3x = 8 - 8
0 = 0
-----------
You solved the 1st equation for y and did that correctly.
Then you substituted that value of y back into the 1st
equation; so you ended up with an identity.
What you need to do is substitute the y-value of the
1st equation into the 2nd equation. Then you will
probably get a correct answer.
Good luck
Cheers,
Stan H.